OpenStudy (logan13):

What is the value of the expression |x| + |y + z| when x = –7, y = 8, and z = –11?

OpenStudy (logan13):

@inowalst

OpenStudy (inowalst):

Plug in the values.

OpenStudy (ankit042):

substitute value in the equation first

OpenStudy (inowalst):

Basically 7+8+11. What is that?

OpenStudy (logan13):

24

OpenStudy (anonymous):

One question. Aren't you supposed to find the absolute value after you solve the question in between the lines?

OpenStudy (inowalst):

Exactly. Absolute value makes negative numbers positive.

OpenStudy (anonymous):

But doesn't it matter if you bake it positive before or after you solve it?

OpenStudy (logan13):

-24

OpenStudy (inowalst):

Not getting you. If you have negative numbers, like two of the numbers we were given, by finding their absolute value, you make them positive. And even if they're already positive, they stay that way.

OpenStudy (inowalst):

@logan13 7+8=15+11=? Not 24.

OpenStudy (logan13):

ohhhh i read it wrong 26

OpenStudy (anonymous):

No I meant that is there an order to if you have to solve it first, then make it absolute value? for example |8+-11| = |-3| = |3|

OpenStudy (inowalst):

You can do that, or just turn them positive. Which basically gives you the same answer.

OpenStudy (logan13):

sooooo 26

OpenStudy (inowalst):

Yes.